OpenStudy (anonymous):

2x+3y=3 x-12y=-9 solve for x and y

OpenStudy (jwt625):

it's wrong

OpenStudy (jwt625):

x=1/3,y=7/9

OpenStudy (jwt625):

the answer should be x=1/3,y=7/9

OpenStudy (anonymous):

the mistake was leaving a -9 instead of a 3 in the substitution of the first equation...

OpenStudy (anonymous):

2x+3y=3 x-12y=-9 make x the subject in the 2nd eqn x=12y-9 put it into the first eqn 2(12y-9)+3y=-9 24y-18+3y=-9 27y-18=-9 27y=-27 y=-1 put y=-1 into 2x+3y=3 2x-3=3 2x=6 x=3 x=3, y=-1